susan is now two thrids pams age. three years ago she was half pams age. six years from now she'll be four fifths pams age. How old arre susan andpam now? write the equation show all work
Answer by stanbon(75887)

susan is now two thirds pams age.
three years ago she was half pams age.
six years from now she'll be four fifths pams age. How old are susan and pam now? write the equation show all work
---------
Equations:
s = (2/3)p
-----
(s-3)= (1/2)(p-3)
-----
(s+6) = (4/5)(p+6)
-------
Note: It is unusual to have three equation statements
for a problem with 2 variables. There may be a contradiction
involved.
=======================
Using equation #1 and #2 you get the following:
Substitute for "s" and solve for "p":
(2/3)p-3 = (1/2)(p-3)
(2/3)p - 2 = (1/2)p - (3/2)
------------------------------
[(2/3)-(1/2)]p = 2-(3/2)
[(4-3)/6]p = (1/2)
(1/6)p = (1/2)
p = 3 yrs old (Pam's age now)
---
Solve for "s":
s = (2/3)p
----
s = (2/3)3
---
s = 2 yrs old
--------------------
Comment: According to that, three years ago Sue would have been -1.
Hmmmmmm????
--------------------
Using equations #2 and #3 you get:
(s-3)= (1/2)(p-3)
(s+6) = (4/5)(p+6)
--------------------------------
2s-6 = p-3
5s+30 = 4p+24
------
2s-p = 3
s -4p = -6
----
2s-p = 3
2s-8p = -12
----
Subtract and solve for "p":
7p = 15
p = 2 1/7 years old (Pam's age now)
---
Solve for "s":
s - 4p = -6
s -4(15/7) = -6
s -60/7 = -42/7
s = 18/7
s = 2 4/7 years old (Sue's age now)
----
Comment:
The equations are incosistent with one another.
====================================================
cheers,
Stan H.
-----------